# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 5 A stable equivalence of Morita type
Since the rank of $V$ is coprime to $p$, by [16, Proposition 5.1.18] and Corollary 3.4, $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. Clearly, the $\mathcal{O}$-rank of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ equals $\mathrm{rank}_{\mathcal{O}}(S^{\circ})\cdot\mathrm{rank}_{\mathcal{O}}(\mathbb{A}_{\delta})$. By Corollary 4.7, $\frac{\mathrm{rank}_{\mathcal{O}}(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})}{|D|}$ is coprime to $p$. By Lemma 5.11 and argument in the paragraph 5.12. applying Lemma 3.3 to $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$, we can get the statements (i) and (ii).
For any nontrivial subgroup $Q$ of $D$, by [22, Proposition 5.6], there is a $D$-interior algebra isomorphism $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)\cong S^{\circ}(Q)\mathop{\otimes}\limits_{k}\mathbb{A}_{\delta}(Q)$, which is isomorphic to $(S(Q))^{\circ}\mathop{\otimes}\limits_{k}\mathbb{A}_{\delta}(Q)$. By Proposition 4.11, $\mathbb{A}_{\delta}(Q)$ is isomorphic to $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Note that $S(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras. Therefore, $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)$ is isomorphic to $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Since $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ has a finite $D\times D$-stable $\mathcal{O}$-basis by Lemma 5.11, we have $((S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q))(D)\cong(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(D)$. This implies that $\mathrm{Br}_{Q}(e)$ belongs to a local point of $D$ on $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)$, which is isomorphic to $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$. Then the statement (iii) follows from [17, Proposition 7.3.10 (iv)]. ∎
@@
Assume that the block $b$ is a hyperfocal abelian Frobenius block. Then there is a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$. Consequently, there is a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$.
@@
By [17, Proposition 6.15.2], the structural map $\mbox{$\mathcal{O}$}L\longrightarrow\mathbb{A}^{\prime}$ splits as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules. We identify $L$ with a subgroup of $(\mathbb{A}^{\prime})^{*}$. Therefore, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{A}^{\prime}=\mbox{$\mathcal{O}$}L\oplus\mathbb{Y}$ for some $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-subbimodules $\mathbb{Y}$ of $\mathbb{A}^{\prime}$. In particular, $\mathbb{Y}$ is a direct summand of $\mathbb{A}^{\prime}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. It is easy to check that $C_{L}(Q)=D$ for any nontrivial subgroup $Q$ of $D$. Then $(\mbox{$\mathcal{O}$}L)(Q)\cong kD$ for any nontrivial subgroup $Q$ of $D$. By Proposition 5.13 (iii), we have $\mathbb{Y}(Q)=0$ for any nontrivial subgroup $Q$ of $D$.
@@
On the other hand, as $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mathbb{A}^{\prime}$ is a direct summand of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$. Hence, $\mathbb{Y}$ is also a direct summand of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the $\mbox{$\mathcal{O}$}(D\times D)$-isomorphism (5.11.1), $\mathbb{Y}$ has to be isomorphic to a direct summand of $\mathbb{X}$. Therefore, every direct summand of $\mathbb{Y}$ has the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. Suppose that $\mathbb{Y}$ has a direct summand $U$ isomorphic to $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ with proper subgroup $1\neq Q$ of $D$. Then there is an element $y^{\prime}_{U}$ of $\mathbb{Y}$ such that for any $(w_{1},w_{2})\in D\times D$, $w_{1}y^{\prime}_{U}w_{2}^{-1}=y^{\prime}_{U}$ if and only if $w_{1}\in Q$ and $w_{2}=hw_{1}h^{-1}$ and $U=\mbox{$\mathcal{O}$}Dy^{\prime}_{U}D$. Then we have for any $(w_{1},w_{2})\in D\times D$, $w_{1}y^{\prime}_{U}\tilde{h}w_{2}^{-1}=y^{\prime}_{U}\tilde{h}$ if and only if $w_{1}\in Q$ and $w_{1}=w_{2}$ and then $\mbox{$\mathcal{O}$}Dy^{\prime}_{U}\tilde{h}D$ is isomorphic to $\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. Since $\mathbb{Y}$ is an $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-subbimodule of $\mathbb{A}^{\prime}$, $\mbox{$\mathcal{O}$}Dy^{\prime}_{U}\tilde{h}D$ is contained in $\mathbb{Y}$ and also a direct summand of $\mathbb{Y}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. But it is easy to check that $(\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$}))(Q)\neq 0$. This is impossible since $\mathbb{Y}(Q)=0$ when $Q\neq 1$. In particular, as an $\mbox{$\mathcal{O}$}(D\times D)$-module, $\mathbb{Y}$ is projective.
@@
Clearly, $\mbox{$\mathcal{O}$}L$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. So as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mbox{$\mathcal{O}$}L$ is isomorphic to a direct summand of $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}D}\mbox{$\mathcal{O}$}L$. Obviously, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{Y}$ is isomorphic to $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}L}\mathbb{Y}\mathop{\otimes}\limits_{\mathcal{O}L}\mbox{$\mathcal{O}$}L$. Therefore, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{Y}$ is isomorphic to a direct summand of $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}D}\mathbb{Y}\mathop{\otimes}\limits_{\mathcal{O}D}\mbox{$\mathcal{O}$}L$. Since $\mbox{$\mathcal{O}$}L$ is a projective $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$-bimodule, $\mbox{$\mathcal{O}$}L$ is also a projective $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodule. Denote the $\mathbb{A}^{\prime}$-$\mbox{$\mathcal{O}$}L$ bimodule $\mathbb{A}^{\prime}$ by $\mathbb{A}^{\prime}_{\mathcal{O}L}$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}^{\prime}$-bimodule $\mathbb{A}^{\prime}$ by ${{}_{\mathcal{O}L}}\mathbb{A}^{\prime}$. Then by [16, Proposition 4.14.12], the bimodules $\mathbb{A}^{\prime}_{\mathcal{O}L}$ and ${{}_{\mathcal{O}L}}\mathbb{A}^{\prime}$ induce a stable equivalence of Mortia type between $\mathbb{A}^{\prime}$ and $\mbox{$\mathcal{O}$}L$. By Proposition 5.13 (ii), the $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $(V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}_{\delta}$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})$ induce a stable equivalence of Mortia type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$. Moreover, by Corollary 3.4, the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a stable equivalence of Mortia type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. ∎
@@
This can be deduced easily from [16, Proposition 4.14.3] and the fact that the Cartan matrix of $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{hyp}})$ has determinant $|D|$. ∎
@@
Keep the notation as above. Denote the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$ by $M_{\mathrm{st}}$. It is clear that $M_{\mathrm{st}}$ is an indecomposable $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule. In order to get the information about source algebras, we need to consider the $\Delta(P)$-equivariant property of the bimodule $M_{\mathrm{st}}$. Here, $\Delta(P)=\{(u,u)\mid u\in P\}$ is the diagonal subgroup of $P\times P$. The following is an investigation of this issue.
@@
Borrowing the notation from Proposition 4.12, there is a unitary subalgebra of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}L$ as $D$-interior algebras, which is still denoted by $\mbox{$\mathcal{O}$}L$ for simplicity. At the same time, there is an element $a_{u,\delta}$ in $N_{A_{\delta}}(D)$ such that $a_{u,\delta}(vj)a_{u,\delta}^{-1}=(uvu^{-1})j$ and $\mathbb{A}_{\delta}^{a_{u,\delta}}=\mathbb{A}_{\delta}$ for any $u\in P$ and any $v\in D$. In particular, the subgroup $j+J(\mathbb{A}_{\delta})^{D}$ is stable under the conjugation action of $a_{u,\delta}$ for any $u\in P$. Recall from the proof of Proposition 4.12 that we have the following split short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\mathbb{A}_{\delta}}(D)/(j+J(\mathbb{A}_{\delta})^{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
Then we have a complement $\hat{E}_{\mathfrak{h}}/(j+J(\mathbb{A}_{\delta})^{D})$ of $k^{*}$ in $N_{\mathbb{A}_{\delta}}(D)/(j+J(\mathbb{A}_{\delta})^{D})$. Denote by $\pi$ the sujective homomorphism of the above sequence, which induces the following another split short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${j+J(\mathbb{A}_{\delta})^{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{E}_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
For simplicity, we still denote by $E_{\mathfrak{h}}$ and $\pi$ a complement of $j+J(\mathbb{A}_{\delta})^{D}$ in $\hat{E}_{\mathfrak{h}}$ and the surjective homomorphism of this sequence, respectively.
@@
Let $u$ be an element of $P$ such that it acts trivially on $E_{\mathfrak{h}}$ through the canonical map $P\longrightarrow P/C_{P}(D)$. Then the element $a_{u,\delta}$ can be chosen such that it acts trivially on $E_{\mathfrak{h}}$ through the conjugation action.

### v2

@@ 5 A stable equivalence of Morita type
Since the rank of $V$ is coprime to $p$, by [15, Proposition 5.1.18] and Corollary 3.4, $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. Clearly, the $\mathcal{O}$-rank of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ equals $\mathrm{rank}_{\mathcal{O}}(S^{\circ})\cdot\mathrm{rank}_{\mathcal{O}}(\mathbb{A}_{\delta})$. By Corollary 4.7, $\frac{\mathrm{rank}_{\mathcal{O}}(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})}{|D|}$ is coprime to $p$. By Lemma 5.11 and the argument in the paragraph 5.12, applying Lemma 3.3 to $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$, we can get the statements (i) and (ii).
For any nontrivial subgroup $Q$ of $D$, by [22, Proposition 5.6], there is a $D$-interior algebra isomorphism $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)\cong S^{\circ}(Q)\mathop{\otimes}\limits_{k}\mathbb{A}_{\delta}(Q)$, which is isomorphic to $(S(Q))^{\circ}\mathop{\otimes}\limits_{k}\mathbb{A}_{\delta}(Q)$. By Proposition 4.13, $\mathbb{A}_{\delta}(Q)$ is isomorphic to $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Note that $S(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras. Therefore, $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)$ is isomorphic to $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Since $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ has a finite $D\times D$-stable $\mathcal{O}$-basis by Lemma 5.11, we have $((S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q))(D)\cong(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(D)$. This implies that $\mathrm{Br}_{Q}(e)$ belongs to a local point of $D$ on $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)$, which is isomorphic to $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$. Then the statement (iii) follows from [16, Proposition 7.3.10 (iv)]. ∎
@@
Assume that the block $b$ is a hyperfocal abelian Frobenius block. Then there is a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$. Consequently, there is a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$.
@@
By [16, Proposition 6.15.2], the structural map $\mbox{$\mathcal{O}$}L\longrightarrow\mathbb{A}^{\prime}$ splits as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules. We identify $L$ with a subgroup of $(\mathbb{A}^{\prime})^{\times}$. Therefore, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{A}^{\prime}=\mbox{$\mathcal{O}$}L\oplus\mathbb{Y}$ for some $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-subbimodules $\mathbb{Y}$ of $\mathbb{A}^{\prime}$. In particular, $\mathbb{Y}$ is a direct summand of $\mathbb{A}^{\prime}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. It is easy to check that $C_{L}(Q)=D$ for any nontrivial subgroup $Q$ of $D$. Then $(\mbox{$\mathcal{O}$}L)(Q)\cong kD$ for any nontrivial subgroup $Q$ of $D$. By Proposition 5.13 (iii), we have $\mathbb{Y}(Q)=0$ for any nontrivial subgroup $Q$ of $D$.
@@
On the other hand, as $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mathbb{A}^{\prime}$ is a direct summand of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$. Hence, $\mathbb{Y}$ is also a direct summand of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the $\mbox{$\mathcal{O}$}(D\times D)$-module decomposition (5.11.1), $\mathbb{Y}$ has to be isomorphic to a direct summand of $\mathbb{X}$ since $\mbox{$\mathcal{O}$}L$ is isomorphic to $\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}Dg$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Therefore, every direct summand of $\mathbb{Y}$ has the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. Suppose that $\mathbb{Y}$ has a direct summand $U$ isomorphic to $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ with a nontrivial proper subgroup $Q$ of $D$. Then there is an element $y^{\prime}_{U}$ in $\mathbb{Y}$ such that for any $(w_{1},w_{2})\in D\times D$, $w_{1}y^{\prime}_{U}w_{2}^{-1}=y^{\prime}_{U}$ if and only if $w_{1}\in Q$ and $w_{2}=h^{-1}w_{1}h$ and $U=\mbox{$\mathcal{O}$}Dy^{\prime}_{U}D$. Then we have for any $(w_{1},w_{2})\in D\times D$, $w_{1}y^{\prime}_{U}\tilde{h}^{-1}w_{2}^{-1}=y^{\prime}_{U}\tilde{h}^{-1}$ if and only if $w_{1}\in Q$ and $w_{1}=w_{2}$ and then $\mbox{$\mathcal{O}$}Dy^{\prime}_{U}\tilde{h}D$ is isomorphic to $\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. Since $\mathbb{Y}$ is an $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-subbimodule of $\mathbb{A}^{\prime}$, $\mbox{$\mathcal{O}$}Dy^{\prime}_{U}\tilde{h}D$ is contained in $\mathbb{Y}$ and also a direct summand of $\mathbb{Y}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. But it is easy to check that $(\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$}))(Q)\neq 0$. This is impossible since $\mathbb{Y}(Q)=0$ when $Q$ is nontrivial. In particular, as an $\mbox{$\mathcal{O}$}(D\times D)$-module, $\mathbb{Y}$ is projective.
@@
Clearly, $\mbox{$\mathcal{O}$}L$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. So $\mbox{$\mathcal{O}$}L$ is isomorphic to a direct summand of $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}D}\mbox{$\mathcal{O}$}L$ as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules. Obviously, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{Y}$ is isomorphic to $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}L}\mathbb{Y}\mathop{\otimes}\limits_{\mathcal{O}L}\mbox{$\mathcal{O}$}L$. Therefore, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{Y}$ is isomorphic to a direct summand of $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}D}\mathbb{Y}\mathop{\otimes}\limits_{\mathcal{O}D}\mbox{$\mathcal{O}$}L$. Since $\mathbb{Y}$ is a projective $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$-bimodule, $\mathbb{Y}$ is also a projective $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodule. Denote the $\mathbb{A}^{\prime}$-$\mbox{$\mathcal{O}$}L$ bimodule $\mathbb{A}^{\prime}$ by $\mathbb{A}^{\prime}_{\mathcal{O}L}$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}^{\prime}$-bimodule $\mathbb{A}^{\prime}$ by ${{}_{\mathcal{O}L}}\mathbb{A}^{\prime}$. Then by [15, Proposition 4.14.12], the bimodules $\mathbb{A}^{\prime}_{\mathcal{O}L}$ and ${{}_{\mathcal{O}L}}\mathbb{A}^{\prime}$ induce a stable equivalence of Mortia type between $\mathbb{A}^{\prime}$ and $\mbox{$\mathcal{O}$}L$. By Proposition 5.13 (ii), the $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $(V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e\cong((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)\mathop{\otimes}\limits_{\mathbb{A}^{\prime}}\mathbb{A}^{\prime}_{\mathcal{O}L}$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}_{\delta}$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cong{{}_{\mathcal{O}L}}\mathbb{A}^{\prime}\mathop{\otimes}\limits_{\mathbb{A}^{\prime}}(e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}))$ induce a stable equivalence of Mortia type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$. Moreover, by Corollary 3.4, the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a stable equivalence of Mortia type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. ∎
@@
This can be deduced easily from [15, Proposition 4.14.13] and the fact that the Cartan matrix of $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ has determinant $|D|$. ∎
@@
Keep the notation as above. Denote the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$ by $M_{\mathrm{st}}$. Then the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $M_{\mathrm{st}}$ is indecomposable. Indeed, it suffices to show that the $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $(V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e$ is indecomposable since the indecomposabilitiy is preserved under a Mortia equivalence. It is easy to check that the endomorphism algebra of the $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $(V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e$ is isomorphic to $(\mathbb{A}^{\prime})^{L}$, which is a local algebra. Since $P$ can act on $E_{\mathfrak{h}}$ through the canonical homomorphism $P\longrightarrow P/C_{P}(D)$, it makes sense to define the twisted $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-module ${{}_{(u,u)}}M_{{\rm st}}$ of $M_{{\rm st}}$ (see the paragraph 4.9) and then consider the $\Delta(P)$-equivariant property of the bimodule $M_{\mathrm{st}}$ which is a key property to get the information about source algebras by Clifford Theory. Here, $\Delta(P)=\{(u,u)\mid u\in P\}$ is the diagonal subgroup of $P\times P$ and the $\Delta(P)$-equivariant property means that $M_{{\rm st}}$ is isomorphic to ${{}_{(u,u)}}M_{{\rm st}}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule for any $(u,u)\in\Delta(P)$. The following is an investigation of this issue.
@@
Borrowing the notation from Proposition 4.16, there is a unitary subalgebra of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}L$ as $D$-interior algebras. More explicitly, there is a subgroup $\{\mathfrak{e}_{\tilde{g}}\mid\tilde{g}\in E_{\mathfrak{h}}\}$ of $N_{\mathbb{A}_{\delta}}(D)$ such that it intersects with $(\mathbb{A}_{\delta}^{D})^{\times}$ trivially and sending $\mathfrak{e}_{\tilde{g}}$ to $\tilde{g}$ gives an isomorphism between it and $E_{\mathfrak{h}}$, which induces the same conjugation actions on $D$. We denote this subgroup of $N_{\mathbb{A}_{\delta}}(D)$ by $\mathfrak{E}_{\mathfrak{h}}$. At the same time, there is an element $a_{u,\delta}$ in $N_{A_{\delta}}(D)$ such that $a_{u,\delta}(vj)a_{u,\delta}^{-1}=(uvu^{-1})j$ and $\mathbb{A}_{\delta}^{a_{u,\delta}}=\mathbb{A}_{\delta}$ for any $u\in P$ and any $v\in D$. Since the $L$-interior algebra structure of $\mathbb{A}_{\delta}$ is essentially given by the group isomorphism between $\mathfrak{E}_{\mathfrak{h}}$ and $E_{\mathfrak{h}}$, the element $a_{u,\delta}$ should be chosen carefully such that $\mathfrak{E}_{\mathfrak{h}}$ is stabilized under its conjugation action on $\mathbb{A}_{\delta}$ and then the induced action on $\mathfrak{E}_{\mathfrak{h}}$ is compatible with the conjugation action of $u$ on $E_{\mathfrak{h}}$ for any $u\in P$. Explicitly, we have the following lemma.
@@
For any $u\in P$, we can choose an element $a_{u,\delta}$ in $N_{A_{\delta}}(D)$ such that ${{}^{a_{u,\delta}}}\mathfrak{E}_{\mathfrak{h}}=\mathfrak{E}_{\mathfrak{h}}$. Furthermore, for any $\tilde{g}\in E_{\mathfrak{h}}$, ${{}^{a_{u,\delta}}}\mathfrak{e}_{\tilde{g}}=\mathfrak{e}_{{{}^{u}}\tilde{g}}$. In particular, when the conjugation action of $u$ on $E_{\mathfrak{h}}$ induced by the canonical map $P\longrightarrow P/C_{P}(D)$ is trivial, we have ${{}^{a_{u,\delta}}}\mathfrak{e}_{\tilde{g}}=\mathfrak{e}_{\tilde{g}}$ for any $\tilde{g}\in E_{\mathfrak{h}}$.